# FocalCodec: Low-Bitrate Speech Coding via Focal Modulation Networks

### v1

@@ 4.5 Downstream Tasks
We evaluate performance on automatic speech recognition (ASR), speaker identification (SI), and speech emotion recognition (SER). These tasks allow us to assess token quality along three axes: semantic information retention (ASR), acoustic information retention (SI), and emotion information retention (SER, which requires a non-trivial combination of semantic and acoustic clues). To focus on the disentanglement of learned representations, we employ shallow downstream models, aiming to stay as close as possible to linear probing. Following (Zhang et al., 2024), we employ a shallow BiLSTM for all tasks. For ASR, we use LibriSpeech (Panayotov et al., 2015) train-clean-100 and train-clean-360 for training, dev-clean for validation, and test-clean for testing. The word error rate (WER) is reported. For SI, we also use LibriSpeech, grouping utterances from train-clean-100 and train-clean-360 by speaker ID. Data are randomly split into training, validation and test sets in a ratio of 80% / 10% / 10%. The speaker error rate (ER) is reported. For SER, we use the IEMOCAP dataset (Busso et al., 2008), focusing on four emotions: sadness, happiness, anger, and neutral. Sessions 1–4 are used for training, session 5F for validation, and session 5M for testing. The emotion ER is reported. Details about the model architecture, hyperparameters, and training procedure are provided in Section D.2.
Discussion
Table 5 shows the results. In ASR, FocalCodec@50 achieves the third lowest WER. While SpeechTokenizer and Stable Codec perform slightly better, the former operates at $\sim$1.5x higher bitrate using two codebooks, while the latter was fine-tuned on force-aligned phoneme data to enhance semantic representations. In contrast, our model is purely self-supervised. In SI, FocalCodec@50 achieves a marginally higher error rate ($\sim$2%) than codecs such as BigCodec and WavTokenizer. However, these models perform significantly worse in ASR due to being trained solely with reconstruction-based objectives. On the other hand, the purely semantic WavLM-KM6 codec performs competitively in ASR but exhibits the highest ER in SI despite using the same encoder as FocalCodec. This further confirms the effectiveness of our codec design, as it improves WER over WavLM-KM6 while preserving speaker information. Interestingly, Stable Codec also performs poorly in SI, likely because semantic fine-tuning tends to remove acoustic information from the representations. In SER, no codec clearly excels, with FocalCodec@50 performing on par with the best models. Overall, FocalCodec@50 shows competitive performance across all discriminative tasks, rivaling hybrid codecs with more complex multi-codebook designs and higher bitrates. The more compressed variants, FocalCodec@25 and FocalCodec@12.5, still achieve acceptable performance while operating at ultra-low bitrates.
@@ Discussion
We evaluate performance on speech enhancement (SE), speech separation (SS), and text-to-speech (TTS). For these tasks, we employ more powerful transformer-based downstream models, focusing on generation quality. For SE we use VoiceBank (Valentini-Botinhao et al., 2016). To form a validation set, we randomly select two speakers from the training set. The input tokens are extracted from noisy utterances, while the target tokens come from clean utterances. Performance metrics include DNSMOS, dWER, and Sim. For SS, we use Libri2Mix (Cosentino et al., 2020) train-100, dev, and test sets. The setup mirrors that of speech enhancement: input tokens are derived from speech mixtures, while target tokens correspond to the two individual sources. For TTS, we use LibriSpeech train-clean-100 and train-clean-360 for training, dev-clean for validation, and test-clean for testing. The input consists of character-based text tokens, while the target tokens are derived from the corresponding utterances. Performance is evaluated using UTMOS, dWER, and Sim. Details about the model architecture, hyperparameters, and training procedure are provided in Section D.2.
Discussion
@@
From Table 5, we observe that in SE, FocalCodec@50 significantly outperforms all other baselines in terms of dWER. A similar trend is observed for SS, where FocalCodec@50 is consistently superior to the other baselines. However, the absolute performance is still far from practical utility, likely due to the loss of information crucial for SS during quantization. Additional results can be found in Table 10. As with discriminative tasks, FocalCodec@25 and FocalCodec@12.5 show degraded performance, likely due to their ultra-low bitrates. However, this trend is reversed for TTS, with FocalCodec@12.5, followed closely by FocalCodec@25, achieving the best overall results. This can be attributed to the fact that, in autoregressive modeling, shorter sequences reduce the computational burden and simplify the task of predicting the next token. FocalCodec@12.5, operating at a frame rate close to that of text with a single codebook, makes next-token prediction easier and more computationally efficient than other methods. This highlights the importance of having compact, powerful, discrete representations for downstream tasks. It is important to note, however, that we trained on only 460 hours of speech, which explains why TTS performance is not state-of-the-art.
@@ 4.6 Ablation Studies
Due to limited computational resources, we perform ablation studies on a smaller variant of FocalCodec. This variant is similar to the 50 Hz model, with the main difference being the model size, as detailed in Section D.1. We focus on the speech resynthesis task. The results are shown in Table 6. Replacing Vocos with HiFi-GAN significantly degrades UTMOS, despite achieving slightly better speaker fidelity. Also note that Vocos offers from 2x to 3x faster inference. Replacing BSQ with vanilla LFQ results in the degradation of dWER and Sim despite similar UTMOS scores. Replacing Snake activations with leaky ReLU causes only a minor degradation in performance. The most significant performance drop occurs when the focal modulation blocks are replaced with Conformer (Gulati et al., 2020) blocks, anti-aliased multi-periodicity (AMP) (Lee et al., 2023) blocks, or linear layers, in this order. This leads to a notable decrease in both dWER and Sim. This analysis further validates our design choices, highlighting the importance of the selected components for achieving optimal performance.
Table 6: Ablation studies.
Compression/ Decompression Block Down/Upscale Activation Quantizer Decoder UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$
Focal modulation Snake BSQ Vocos 4.14 2.54 95.3
Focal modulation Snake BSQ HiFi-GAN 3.73 2.54 95.7
Focal modulation Snake LFQ HiFi-GAN 3.74 2.75 95.4
Focal modulation Leaky ReLU LFQ HiFi-GAN 3.72 2.85 95.2
Conformer Snake LFQ HiFi-GAN 3.74 3.58 94.3
AMP Snake LFQ HiFi-GAN 3.70 4.52 94.3
Linear Snake LFQ HiFi-GAN 2.55 9.37 82.5
@@ 5 Conclusions
In this work, we introduced FocalCodec, a low-bitrate single-codebook speech codec that employs a novel architecture based on focal modulation. It delivers competitive performance in speech resynthesis and voice conversion at low and ultra-low bitrates while maintaining robustness across diverse conditions, including multilingual and noisy speech. Furthermore, FocalCodec effectively preserves both semantic and acoustic information and provides powerful discrete representations for generative downstream tasks.
@@ Appendix A Limitations
The proposed codec comes with some inherent trade-offs. From an architectural perspective, it is not causal. While it can be streamed in a chunk-wise fashion (see Section F.3), its latency remains too high for real-time applications. Future work should explore alternative designs and/or training strategies to enable a fully causal architecture. Regarding the dataset, aside from the 94,000 hours of pretraining data used for the WavLM encoder, our compressor, quantizer, decompressor, and decoder were trained only on a few hundred hours of clean English speech sampled at 16 kHz. Expanding the dataset to include more data, a broader range of domains (e.g. multilingual speech, mixtures, etc.) and increasing the sampling rate (e.g. 24 kHz) could further improve quality, robustness, and versatility of the model. Additionally, the model is limited to speech and does not support high-fidelity reconstruction of music or environmental sounds.
Appendix B Datasets
@@
  • LibriSpeech (Panayotov et al., 2015) is a large-scale corpus of English read speech derived from audiobooks in the LibriVox project. It contains approximately 1000 hours of speech sampled at 16 kHz, with predefined training, validation, and test splits.
@@
  • LibriTTS (Zen et al., 2019) is a corpus designed for text-to-speech research, constructed from the same source as LibriSpeech. It consists of 585 hours of transcribed speech with predefined training, validation, and test splits.
@@
  • Multilingual LibriSpeech (Pratap et al., 2020) is an extension of LibriSpeech to multiple languages, including English, German, Dutch, French, Spanish, Italian, Portuguese and Polish. It provides approximately 44,500 hours of transcribed English speech and about 6000 hours from other languages.
@@
  • VoiceBank (Valentini-Botinhao et al., 2016) is a dataset primarily used for speech enhancement, including 11,572 utterances from 28 speakers in the training set (noise at 0 dB, 5 dB, 10 dB, and 15 dB), and 872 utterances from 2 unseen speakers in the test set (noise at 2.5 dB, 7.5 dB, 12.5 dB, and 17.5 dB).
@@
  • LibriMix (Cosentino et al., 2020) is a dataset for speech separation and enhancement, created by mixing LibriSpeech utterances with noise from the WHAM! (Wichern et al., 2019) corpus. It provides mixtures of two or three speakers at different signal-to-noise ratios.
@@
  • VCTK (Yamagishi et al., 2017) is a corpus of English speech recordings from 110 speakers with various accents. It is widely used for speaker adaptation, text-to-speech, and voice conversion tasks.
@@
  • IEMOCAP (Busso et al., 2008) is a dataset designed for emotion recognition, consisting of scripted and improvised dialogues performed by 10 actors. It includes audio, video, and textual transcriptions with emotion labels such as happiness, sadness, and anger.
Appendix C Baselines
@@
Additional information about the baseline codecs is provided in Table 7. For our WavLM6-KM (Wang et al., 2024) reproduction, we use LibriSpeech train-clean-100 and train-clean-360. First, we train a k-means quantizer with 512 centroids on top of layer-6 representations from WavLM-large. We train on audio chunks of 16,000 samples with a large batch size of 512 for improved stability, and we stop training when cluster centroids stop changing significantly. Then, we train a dequantizer to minimize the L2 loss between quantized and original WavLM features. We employ a Conformer (Gulati et al., 2020) encoder with 6 layers, 4 attention heads, a hidden dimension of 512, and a feed-forward layer dimension of 512. We train on audio chunks of 7040 samples with a batch size of 16. We use the AdamW (Loshchilov & Hutter, 2019) optimizer with an initial learning rate of 0.0005, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, weight decay of 0.01, and dropout of 0.1. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 5. Training stops when validation loss does not decrease for several consecutive epochs. Finally, we train a HiFi-GAN V1 (Kong et al., 2020) decoder on audio chunks of 7040 samples with a batch size of 16. We use the AdamW optimizer with an initial learning rate of 0.0002, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate follows an exponential decay schedule with a factor of 0.999. Training continues until perceived audio quality stops improving.
@@
Codec Causal Training Datasets Hours Multilingual Audio Domain Checkpoint
EnCodec (Défossez et al., 2023) Optional DNS, CommonVoice, AudioSet, FSD50K, Jamendo 17,000+ Yes General encodec_24khz
DAC (Kumar et al., 2023) No DAPS, DNS, CommonVoice, VCTK, MUSDB, Jamendo 10,000+ Yes General weights_16khz.pth
@@
WavLM6-KM (Wang et al., 2024) No Subset of LibriSpeech (in addition to Libri-Light, GigaSpeech, and VoxPopuli English for WavLM pretraining) 460 (+ 94,000) No Speech discrete-wavlm-codec
SpeechTokenizer (Zhang et al., 2024) No LibriSpeech 960 No Speech speechtokenizer_hubert_avg
@@
SemantiCodec (Liu et al., 2024) No GigaSpeech, subset of OpenSLR, Million Song Dataset, MedleyDB, MUSDB18, AudioSet, WavCaps, VGGSound 20,000+ Yes General semanticodec_tokenrate_50
@@
Mimi (Défossez et al., 2024) Yes Predominantly English speech (in addition to Libri-Light, GigaSpeech, and VoxPopuli English for WavLM pretraining) 7,000,000 (+ 94,000) Likely Speech mimi
@@
WavTokenizer (Ji et al., 2024) No LibriTTS, VCTK, subset of CommonVoice, subset of AudioSet, Jamendo, MUSDB 8000 Yes General WavTokenizer-large-unify-40token
BigCodec (Xin et al., 2024) No LibriSpeech 960 No Speech bigcodec.pt
@@
Stable Codec (Parker et al., 2024) Optional Libri-Light, Multilingual LibriSpeech English 105,000 No Speech stable-codec-speech-16k
Appendix D Hyperparameters and Training Details
D.1 FocalCodec

### v2

@@ 4.5 Downstream Tasks
We evaluate performance on automatic speech recognition (ASR), speaker identification (SI), and speech emotion recognition (SER). These tasks allow us to assess token quality along three axes: semantic information retention (ASR), acoustic information retention (SI), and emotion information retention (SER, which requires a non-trivial combination of semantic and acoustic clues). To focus on the disentanglement of learned representations, we employ shallow downstream models, aiming to stay as close as possible to linear probing. Following zhang2024speechtokenizer , we employ a shallow BiLSTM for all tasks. For ASR, we use LibriSpeech panayotov2015librispeech train-clean-100 and train-clean-360 for training, dev-clean for validation, and test-clean for testing. The word error rate (WER) is reported. For SI, we also use LibriSpeech, grouping utterances from train-clean-100 and train-clean-360 by speaker ID. Data are randomly split into training, validation and test sets in a ratio of 80% / 10% / 10%. The speaker error rate (ER) is reported. For SER, we use the IEMOCAP dataset busso2008iemocap , focusing on four emotions: sadness, happiness, anger, and neutral. Sessions 1-4 are used for training, session 5F for validation, and session 5M for testing. The emotion ER is reported. Details about the model architecture, hyperparameters, and training procedure are provided in Section˜E.2.
Table˜4 shows the results. In ASR, FocalCodec@50 achieves the third lowest WER. While SpeechTokenizer and Stable Codec perform slightly better, the former operates at $\sim$1.5x higher bitrate using two codebooks, while the latter was fine-tuned on force-aligned phoneme data to enhance semantic representations. In contrast, our model is purely self-supervised. In SI, FocalCodec@50 achieves a marginally higher error rate ($\sim$2%) than codecs such as BigCodec and WavTokenizer. However, these models perform significantly worse in ASR due to being trained solely with reconstruction-based objectives. On the other hand, the purely semantic WavLM-KM6 codec performs competitively in ASR but exhibits the highest ER in SI despite using the same encoder as FocalCodec. This further confirms the effectiveness of our codec design, as it improves WER over WavLM-KM6 while preserving speaker information. Interestingly, Stable Codec also performs poorly in SI, likely because semantic fine-tuning tends to remove acoustic information from the representations. In SER, no codec clearly excels, with FocalCodec@50 performing on par with the best models. Overall, FocalCodec@50 shows competitive performance across all discriminative tasks, rivaling hybrid codecs with more complex multi-codebook designs and higher bitrates. The more compressed variants, FocalCodec@25 and FocalCodec@12.5, still achieve good performance while operating at ultra-low bitrates.
@@
We evaluate performance on speech enhancement (SE), speech separation (SS), and text-to-speech (TTS). For these tasks, we employ more powerful transformer-based downstream models, focusing on generation quality. For SE we use VoiceBank valentinibotinhao2016voicebank . To form a validation set, we randomly select two speakers from the training set. The input tokens are extracted from noisy utterances, while the target tokens come from clean utterances. Performance metrics include DNSMOS, dWER, and Sim. For SS, we use Libri2Mix cosentino2020librimix train-100, dev, and test sets. The setup mirrors that of speech enhancement: input tokens are derived from speech mixtures, while target tokens correspond to the two individual sources. For TTS, we use LibriSpeech train-clean-100 and train-clean-360 for training, dev-clean for validation, and test-clean for testing. Note that test-clean contains several utterances longer than 20 seconds ($\sim$4%), whereas our training splits include almost none. To reduce the mismatch between training and testing conditions, we removed these long utterances from the test set. The input consists of character-based text tokens, while the target tokens are derived from the corresponding utterances. Performance is evaluated using UTMOS, dWER, and Sim. Details about the model architecture, hyperparameters, and training procedure are provided in Section˜E.2.
@@
From Table˜4, we observe that in SE, FocalCodec@50 significantly outperforms all other baselines in terms of dWER. A similar trend is observed for SS, where FocalCodec@50 is consistently superior to the other baselines. However, the absolute performance is still far from practical utility, likely due to the loss of information crucial for SS during quantization. As with discriminative tasks, FocalCodec@25 and FocalCodec@12.5 show degraded performance, due to their ultra-low bitrates. However, this trend is reversed for TTS, with FocalCodec@25 achieving the best overall results, followed closely by FocalCodec@12.5. This can be attributed to the fact that, in autoregressive modeling, shorter sequences reduce the computational burden and simplify the task of predicting the next token. Both models, operating at a frame rate closer to that of text with a single codebook, make next-token prediction easier and more computationally efficient than other methods. This highlights the importance of having compact representations for downstream tasks. Note, however, that we trained on only 460 hours of speech, which explains why TTS performance is not state-of-the-art.
Table 4: Evaluation on discriminative and generative downstream tasks.
Codec Bitrate (kbps) $\downarrow$ Discriminative Tasks Generative Tasks
ASR SI SER SE $S$ TTS
Reference 3.56 0.00 100.0 3.77 0.00 100.0 4.09 0.00 100.0
EnCodec 1.50 27.89 3.00 47.00 3.11 37.10 85.9 3.11 78.51 87.3 1.71 64.28 83.2
DAC 1.00 35.89 3.27 45.90 3.03 67.65 81.7 2.76 106.00 83.3 1.34 47.06 85.9
WavLM6-KM 0.45 19.04 22.30 42.90 3.52 22.85 83.6 3.49 76.91 85.0 3.74 38.67 88.7
SpeechTokenizer 1.00 14.97 2.73 41.50 3.21 29.82 85.9 3.13 83.99 87.3 2.69 35.46 89.2
SemantiCodec 0.65 41.42 15.90 51.60 3.59 102.00 83.3 3.59 123.00 84.4 2.82 48.38 91.4
Mimi 0.69 22.98 5.43 44.70 3.30 53.98 84.6 3.41 93.23 88.1 3.11 28.63 93.6
WavTokenizer 0.48 35.62 2.44 49.80 3.41 51.75 88.6 3.54 105.00 86.4 3.68 47.56 92.8
BigCodec 1.04 26.41 2.34 47.50 3.52 26.68 93.2 3.54 89.24 89.4 3.43 54.43 89.4
Stable Codec 0.70 16.85 16.50 46.54 3.55 35.57 82.8 3.61 103.00 78.2 3.19 49.28 88.8
FocalCodec @50 0.65 17.63 4.48 45.60 3.47 10.93 91.4 3.71 73.87 89.0 4.11 28.10 93.3
FocalCodec @25 0.33 21.12 6.07 46.80 3.49 14.74 90.0 3.69 99.96 85.4 4.16 16.75 91.6
FocalCodec @12.5 0.16 33.24 11.69 46.30 3.58 36.98 86.9 3.57 116.00 80.8 4.12 21.59 90.8
@@ 4.6 Ablation Studies
Table 5: Ablation studies on LibriSpeech test-clean panayotov2015librispeech .
Compression
Block
Downscale
Activation
Quantizer
UTMOS $\uparrow$
dWER $\downarrow$
Sim $\uparrow$
Focal modulation Snake BSQ 3.73 2.54 95.7
Focal modulation Snake FSQ 3.71 2.61 94.8
Focal modulation Snake LFQ 3.74 2.75 95.4
Focal modulation Leaky ReLU LFQ 3.72 2.85 95.2
Conformer Snake LFQ 3.74 3.58 94.3
AMP Snake LFQ 3.70 4.52 94.3
Linear Snake LFQ 2.55 9.37 82.5
Due to limited computational resources, we perform ablation studies on a smaller variant of FocalCodec. This variant is similar to the 50 Hz model, with the main difference being the model size, as detailed in Section˜E.1. We focus on the clean speech resynthesis task using LibriSpeech test-clean panayotov2015librispeech . The results are shown in Table˜5. Replacing BSQ with FSQ mentzer2024finite leads to worse UTMOS, dWER, and Sim. It also results in less uniform code usage, as evidenced by the normalized entropy measured for these two configurations (99.7 for BSQ vs. 97.7 for FSQ). Replacing BSQ with vanilla LFQ results in worse dWER and Sim despite similar UTMOS. Replacing Snake activations with leaky ReLU causes only minor performance degradation. The most significant performance drop occurs when the focal modulation blocks are replaced with Conformer gulati2020conformer blocks, anti-aliased multi-periodicity (AMP) lee2023bigvgan blocks, or linear layers, in this order. This leads to a notable decrease in both dWER and Sim. This analysis further validates our design choices, highlighting the importance of the selected components for achieving optimal performance.
@@ 5 Conclusions
In this work, we introduced FocalCodec, a low-bitrate single-codebook speech codec that employs a novel architecture based on focal modulation. It delivers competitive performance in speech resynthesis and voice conversion at low and ultra-low bitrates while maintaining robustness across diverse conditions, including multilingual and noisy speech. Furthermore, it effectively preserves both semantic and acoustic information, providing powerful discrete representations for downstream tasks. A detailed discussion of the limitations and societal impact of this work is provided in Appendices˜A and B.
6 Acknowledgments
We acknowledge support from NSERC, the Digital Research Alliance of Canada (alliancecan.ca), the NVIDIA Academic Grant Program for computing resources, and Translated for funding through the Immediate research grant.
@@ Appendix A Limitations
Despite its competitive performance, FocalCodec is undertrained compared to other state-of-the-art approaches. While the WavLM encoder benefits from 94k hours of pretraining, the rest of the pipeline was trained on only a few hundred hours of clean English speech. Expanding the dataset to include more data, a broader range of domains (e.g. multilingual speech, mixtures, etc.) could further improve quality, robustness, and versatility of the model. By comparison, competing methods such as WavTokenizer (8k hours), StableCodec (105k hours), and Mimi (7M hours) are trained on significantly larger and more diverse datasets.
Appendix B Societal Impact
We believe this research has the potential for meaningful societal benefits. Ultra-low bitrate speech codecs can significantly reduce the bandwidth and storage requirements for transmitting and storing spoken content. This has practical implications for improving the accessibility and efficiency of voice communication in bandwidth-constrained settings, such as rural or remote areas, and for enabling on-device speech applications with minimal resource consumption. However, we also acknowledge potential risks associated with misuse. In particular, voice conversion capabilities enabled by FocalCodec could potentially be exploited for malicious purposes, including voice cloning, impersonation, and the creation of deceptive or harmful deepfake audio content. To mitigate these risks, we encourage responsible use of this technology and further research into detection and authentication mechanisms to ensure secure and ethical deployment. It is worth noting nevertheless that similar capabilities are already publicly accessible, with models like https://huggingface.co/amphion/Vevo offering few-shot voice conversion.
Appendix C Datasets
@@
  • LibriSpeech [47] is a large-scale corpus of English read speech derived from audiobooks in the LibriVox project. It contains approximately 1000 hours of speech sampled at 16 kHz, with predefined training, validation, and test splits. License: CC BY 4.0.
@@
  • LibriTTS [84] is a corpus designed for text-to-speech research, constructed from the same source as LibriSpeech. It consists of 585 hours of transcribed speech with predefined training, validation, and test splits. License: CC BY 4.0.
@@
  • Multilingual LibriSpeech [51] is an extension of LibriSpeech to multiple languages, including English, German, Dutch, French, Spanish, Italian, Portuguese and Polish. It provides approximately 44,500 hours of transcribed English speech and about 6000 hours from other languages. License: CC BY 4.0.
@@
  • VoiceBank [66] is a dataset primarily used for speech enhancement, including 11,572 utterances from 28 speakers in the training set (noise at 0 dB, 5 dB, 10 dB, and 15 dB), and 872 utterances from 2 unseen speakers in the test set (noise at 2.5 dB, 7.5 dB, 12.5 dB, and 17.5 dB). License: CC BY 4.0.
@@
  • LibriMix [12] is a dataset for speech separation and enhancement, created by mixing LibriSpeech utterances with noise from the WHAM! [73] corpus. It provides mixtures of two or three speakers at different signal-to-noise ratios. License: MIT.
@@
  • VCTK [77] is a corpus of English speech recordings from 110 speakers with various accents. It is widely used for speaker adaptation, text-to-speech, and voice conversion tasks. License: CC BY 4.0.
@@
  • IEMOCAP [7] is a dataset designed for emotion recognition, consisting of scripted and improvised dialogues performed by 10 actors. It includes audio, video, and textual transcriptions with emotion labels such as happiness, sadness, and anger. License: https://sail.usc.edu/iemocap/iemocap_release.htm.
Appendix D Baselines
@@
Additional information about the baseline codecs is provided in Table˜6. For our WavLM6-KM [71] reproduction, we use LibriSpeech train-clean-100 and train-clean-360. First, we train a k-means quantizer with 512 centroids on top of layer-6 representations from WavLM-large. We train on audio chunks of 16,000 samples with a large batch size of 512 for improved stability, and we stop training when cluster centroids stop changing significantly. Then, we train a dequantizer to minimize the L2 loss between quantized and original WavLM features. We employ a Conformer [19] encoder with 6 layers, 4 attention heads, a hidden dimension of 512, and a feed-forward layer dimension of 512. We train on audio chunks of 7040 samples with a batch size of 16. We use the AdamW [40] optimizer with an initial learning rate of 0.0005, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, weight decay of 0.01, and dropout of 0.1. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 5. Training stops when validation loss does not decrease for several consecutive epochs. Finally, we train a HiFi-GAN V1 [32] decoder on audio chunks of 7040 samples with a batch size of 16. We use the AdamW optimizer with an initial learning rate of 0.0002, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate follows an exponential decay schedule with a factor of 0.999. Training continues until perceived audio quality stops improving.
@@
Codec Causal Training Datasets Hours Multilingual Audio Domain Checkpoint License
EnCodec [14] Optional DNS, CommonVoice, AudioSet, FSD50K, Jamendo 17k+ Yes General encodec_24khz MIT
DAC [34] No DAPS, DNS, CommonVoice, VCTK, MUSDB, Jamendo 10k+ Yes General weights_16khz.pth MIT
@@
WavLM6-KM [71] No Subset of LibriSpeech (in addition to Libri-Light, GigaSpeech, and VoxPopuli English for WavLM pretraining) 460 (+ 94k) No Speech discrete-wavlm-codec Apache 2.0
SpeechTokenizer [86] No LibriSpeech 960 No Speech speechtokenizer_hubert_avg Apache 2.0
@@
SemantiCodec [37] No GigaSpeech, subset of OpenSLR, Million Song Dataset, MedleyDB, MUSDB18, AudioSet, WavCaps, VGGSound 20k+ Yes General semanticodec_tokenrate_50 MIT
@@
Mimi [16] Yes Predominantly English speech (in addition to Libri-Light, GigaSpeech, and VoxPopuli English for WavLM pretraining) 7M (+ 94k) Likely Speech mimi CC BY 4.0
@@
WavTokenizer [25] No LibriTTS, VCTK, subset of CommonVoice, subset of AudioSet, Jamendo, MUSDB 8k Yes General WavTokenizer-large-unify-40token MIT
BigCodec [76] No LibriSpeech 960 No Speech bigcodec.pt MIT
@@
Stable Codec [48] Optional Libri-Light, Multilingual LibriSpeech English 105k No Speech stable-codec-speech-16k StabilityAI
Appendix E Hyperparameters and Training Details
E.1 FocalCodec